Week of 191104

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday - virtual meeting

  • You may provide relevant incidents, announcements etc. here for the operations record.


  • local:
  • remote:

Experiments round table:

  • ATLAS reports ( raw view) -
    • Normal operations
    • "Let's Encrypt" host certificates were opted to be used by two US sites
      • They are not supported by ATLAS since:
        • They are not included in the IGTF repository
        • They do not support CRL (Certificate Revocation List) i.e.
        • It does not work with the current XRootD configuration
      • What's WLCG stand on the topic? Supported? Not? To be re-discussed next week.

  • ALICE -
    • NTR

Sites / Services round table:

  • ASGC:
    • End of downtime just now, checking the services (it's so far so good now).
  • BNL:
  • CNAF:
  • EGI:
  • FNAL: On October 25 water service to the building housing CMS tape libraries was cut due to a water leak, and humidity levels dropped to a point requiring pausing the tape libraries. By October 29 temporary humidifiers had been deployed and tape libraries unpaused. Fermilab Facilities and Engineering Services expedited work to repair the water leak, with full service restored on the 30th. Net unavailability of tape in this incident about 3.5 days.
  • IN2P3: NTR
  • KISTI:
  • KIT: Resolved the issues surrounding ARC-CEs 3 and 4 of last week this early morning (though confirmation from LHCb is pending).
  • NDGF: dCache upgraded to 6.0.0 last week. A few pools remain at 5.2, but will be upgraded during the week.
  • NL-T1: NTR
  • NRC-KI:
  • OSG:
  • PIC:
  • RAL:

  • CERN computing services:
  • CERN storage services: EOSALICE will be upgraded on Wed Nov 06 starting at 10:00 to a new namespace architecture which implies new head node and a dns change. Reads/writes will be affected during that time frame (~30 mins) OTG0052897
  • CERN databases:
  • Monitoring:
  • MW Officer:
  • Networks:
  • Security: NTR


Edit | Attach | Watch | Print version | History: r14 < r13 < r12 < r11 < r10 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r14 - 2019-11-11 - VictorZhiltsov
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback